Question Two

In the economy of Queensland, two products (Sta and Star) are prominent due to their

quality and durability. At the price of GHȼ 30 for Sta, 5,000 quantities of Sta are demanded.

Stonewall, a banker lives in this economy and earns a weekly income of GHȼ 2,000. At this

income level, Stonewall consumes 100 quantities of Star monthly. The price of Sta

increases by 50%, now 4,000 quantities of Sta are demanded. At the initial price of Sta,

1000 quantities of Star were demanded but at the new price of Sta, 3,000 quantities of Star

are demanded. Stonewall attains a professional qualification thus his weekly income

increases by 20%, at the new income level; he consumes 200 quantities of Star monthly.

a) Determine the Own Price Elasticity of Demand for Sta and interpret your answer?

3 Marks

b) Determine the Cross Price Elasticity of Demand for both products and interpret the

relationship between the two products? 3 Mar


1
Expert's answer
2021-06-06T23:20:03-0400

a)


"\frac{4000-5000}{45-30}\times \frac{45+30}{4000+5000}=-0.56"

b)


"\frac{3000-1000}{45-30}\times \frac{45+30}{3000+1000}=2.5"